### v1

@@ 5.3 Systematic search for superintegrable systems
Considering equations (118) step by step for the values of arbitrary parameters presented in (119)-(122) we find the following solutions:
$F=c_{1}x_{1}^{2}+c_{2},\ G=c_{1}x_{2}^{2}+c_{3}$ (124)
for the case (119),
$F=c_{1}x_{1}^{2}+c_{2}+\frac{c_{3}}{x_{1}^{2}},\ G=c_{4}x_{2}^{2}+c_{5}+\frac{c_{6}}{x_{2}^{2}}$ (125)
for the case (120),
$F=c_{1}+\frac{c_{2}}{x_{1}^{2}},\ F=c_{3}+\frac{c_{4}}{x_{1}^{2}}$ (126)
for the case (121),
$F=c_{1}x_{1}^{2}+c_{2}+\frac{c_{3}}{x_{1}^{2}},\ G=4c_{1}x_{2}^{2}+c_{5}+{c_{6}}{x_{2}}$ (127)
for the case (122), and
$F=\frac{c_{1}}{x_{1}^{2}}+c_{2},\ G=c_{3}+{c_{4}}{x_{2}}$ (128)
@@
Version (130) correspond to the system invariant w.r.t. two parametric Lie group including shifts and rotations around the third coordinate axis and so can be omitted as considered before in in paper [45]. Versions (132), (127), and (133) (128) are presented in Items 3, 4 of Table 4 and Items 2, 4 of Table 5 correspondingly.
We have analysed all possible superintegrable systems admitting integral of motion (90) and find seven versions of such systems. In analogous way it is possible to classify the superintegrable systems admitting integrals of motion (106) and find four more superintegrable systems presented in Items 5, 6, 7 of Table 4 and Item 3 of Table 5. Notice that there are two superintegrable systems admitting integrals of motion (90) and (106) both, see Item 4 of Table 1 and Item 2 of Table 5, and there are no superintegrable systems which do not admit either integral (90) or integral (106).
No $f$ $V$ Integrals of motion
1 $\frac{1}{c_{1}x_{1}+c_{2}x_{2}}$ $\frac{c_{3}x_{1}+c_{4}x_{2}+c_{5}}{c_{1}x_{1}+c_{2}x_{2}}$ $\begin{array}[]{l}P_{1}^{2}-P_{2}^{2}+c_{3}x_{1}-c_{4}x_{2}\\ +\left((c_{2}x_{2}-c_{1}x_{1})\cdot H\right),\\ 2P_{1}P_{2}+c_{3}x_{2}+c_{4}x_{1}\\ -((c_{1}x_{2}+c_{2}x_{1})\cdot H),\\ 2P_{3}P_{1}-(c_{1}x_{3}\cdot H)+c_{3}x_{3},\\ 2P_{3}P_{2}-(c_{2}x_{3}\cdot H)+c_{4}x_{3},\\ \{P_{3},L_{2}\}-\frac{c_{1}}{2}(x_{3}^{2}\cdot H)+\frac{1}{2}c_{3}x_{3}^{2},\\ \{P_{3},L_{1}\}+\frac{c_{2}}{2}(x_{3}^{2}\cdot H)-\frac{1}{2}c_{4}x_{3}^{2}\end{array}$
2 $\frac{x_{1}^{2}x_{2}^{2}}{c_{1}x_{1}^{2}+c_{2}x_{2}^{2}+c_{3}x_{1}^{2}x_{2}^{2}}$ $\frac{c_{4}x_{1}^{2}+c_{5}x_{2}^{2}+c_{6}x_{1}^{2}x_{2}^{2}}{c_{1}x_{1}^{2}+c_{2}x_{2}^{2}+c_{3}x_{1}^{2}x_{2}^{2}}$ $\begin{array}[]{l}L_{3}^{2}-\left(\frac{c_{1}x_{1}^{4}+c_{2}x_{2}^{4}}{x_{1}^{2}x_{2}^{2}}\cdot H\right)+\frac{c_{5}x_{1}^{4}+c_{6}x_{2}^{4}}{x_{1}^{2}x_{2}^{2}},\\ L_{1}^{2}-\left(\frac{c_{1}x_{3}^{2}}{x_{2}^{2}}\cdot H\right)+\frac{c_{4}x_{3}^{2}}{x_{2}^{2}},\\ \{P_{2},L_{1}\}-\left(\frac{2c_{1}x_{3}}{x_{2}^{2}}\cdot H\right)+\frac{2c_{4}x_{3}}{x_{2}^{2}},\\ P_{1}^{2}-P_{2}^{2}+\left(\frac{c_{1}x_{1}^{2}-c_{2}x_{2}^{2}}{x_{1}^{2}x_{2}^{2}}\cdot H\right)-\frac{c_{4}x_{1}^{2}-c_{5}x_{2}^{2}}{x_{1}^{2}x_{2}^{2}},\\ \{P_{3},D\}-2(c_{3}x_{3}\cdot H)+2c_{6}x_{3},\\ D^{2}-\left(c_{3}x_{3}r^{2}\cdot H\right)+c_{6}x_{3}r^{2}\end{array}$
3 $\frac{\tilde{r}x_{1}^{2}}{c_{1}x_{2}+c_{2}\tilde{r}x_{1}^{2}+c_{3}\tilde{r}}$ $\frac{c_{4}x_{2}+c_{5}\tilde{r}x_{1}^{2}+c_{6}\tilde{r}}{c_{1}x_{2}+c_{2}\tilde{r}x_{1}^{2}+c_{3}\tilde{r}}$ $\begin{array}[]{l}L_{3}^{2}-\left(\frac{c_{1}\tilde{r}x_{2}+c_{3}x_{2}^{2}}{x_{1}^{2}}\cdot H\right)-\frac{c_{4}\tilde{r}x_{2}+c_{6}x_{2}^{2}}{x_{1}^{2}},\\ \{P_{1},L_{3}\}+\left(\frac{2c_{3}x_{2}\tilde{r}+c_{1}(x_{1}^{2}+2x_{2}^{2})}{x_{1}^{2}\tilde{r}}\cdot H\right)\\ -\frac{2c_{6}x_{2}\tilde{r}+c_{4}(x_{1}^{2}+2x_{2}^{2})}{x_{1}^{2}\tilde{r}},\\ \{P_{3},D\}-\left(2c_{2}x_{3}\cdot H\right)+2c_{5}x_{3},\\ D^{2}-\left(c_{2}r^{2}\cdot H\right)+c_{5}r^{2},\\ \{L_{2},L_{3}\}-\frac{2c_{6}x_{3}\tilde{r}+c_{4}x_{3}(x_{1}^{2}+2x_{2}^{2})}{x_{1}^{2}\tilde{r}}\\ +\left(\frac{2c_{3}x_{3}\tilde{r}+c_{1}x_{3}(x_{1}^{2}+2x_{2}^{2})}{x_{1}^{2}\tilde{r}}\cdot H\right)\end{array}$
4 $\frac{x_{1}^{2}}{x_{2}x_{1}^{2}+c_{1}}$ $\frac{c_{2}x_{2}x_{1}^{2}+c_{3}+c_{4}x_{1}^{2}}{x_{2}x_{1}^{2}+c_{1}}$ $\begin{array}[]{l}L_{2}^{2}-\left(\frac{c_{1}x_{3}^{2}}{x_{1}^{2}}\cdot H\right)+\frac{c_{3}x_{3}^{2}}{x_{1}^{2}},\\ \{P_{1},L_{2}\}-\left(\frac{2c_{1}x_{3}}{x_{1}^{2}}\cdot H\right)+\frac{2c_{3}x_{3}}{x_{1}^{2}},\\ \{P_{2},D\}-\frac{1}{2}\left((x_{1}^{2}+4x_{2}^{2}+x_{3}^{2})\cdot H\right)\\ +\frac{c_{2}}{2}(x_{1}^{2}+4x_{2}^{2}+x_{3}^{2})+2c_{4}x_{2},\\ P_{1}^{2}-P_{2}^{2}+\left(\frac{x_{2}x_{1}^{2}-c_{1}}{x_{1}^{2}}\cdot H\right)-\frac{c_{2}x_{2}x_{1}^{2}-c_{3}}{x_{1}^{2}},\\ P_{2}P_{3}-(c_{1}x_{3}\cdot H)+c_{3}x_{3},\\ \{P_{3},L_{1}\}+\left(\frac{c_{1}x_{3}^{2}}{2}\cdot H\right)-\frac{1}{2}c_{3}x_{3}^{2}\end{array}$
5.4 Integrals of motion dependent on $x_{3}$

### v3

@@ 4.1.2 Systems admitting integral of motion $Q^{(1)}$
The next step is to find inequivalent solutions of equations (98), (99) for all versions of arbitrary parameters enumerated in the above.
The related equation (99) is reduced to the condition $N_{3}=0$ and there are four versions of equations (98) corresponding to the conditions (109), (110), and (111) .
Let us differentiate the first of equations (98) with respect to $x_{2}$ and the second of these equations with respect to $x_{1}$ and equate the obtained expressions. As a result we come to the differential consequence of (98) which does not include unknown $N$:
$\begin{split}&\left(\sigma(x_{2}^{2}-x_{1}^{2})-2\nu x_{2}+\omega\right)M_{1,2}+(\sigma x_{1}x_{2}-\nu x_{1}-\kappa)(M_{1,1}-M_{2,2})\\ &+3(\sigma x_{2}-\nu)M_{1}-3\sigma x_{1}M_{2}=0.\end{split}$ (112)
If conditions (108) and (109) are valid equations (112) and (98) reduced to the following forms:
$x_{1}(M_{2,2}-M_{1,1})-2x_{2}M_{1,2}-3M_{1}=0$ (113)
and
$\begin{split}&2x_{2}M_{1}-x_{1}M_{2}-N_{1}=0,\\ &x_{1}M_{1}+N_{2}=0.\end{split}$ (114)
It is not too easy task to find the generic solution of the above presented equations. However, changing the dependent and independent variables $(x_{1},x_{2})\to(z_{1},z_{2}),\ M(x_{1},x_{2})\to{\Phi(z_{+},z_{-})}$ where
$\Phi(z_{+},z_{-})=\tilde{r}M(x_{1},x_{2}),\ z_{\pm}=\tilde{r}\pm x_{2}$ (115)
it is possible to reduce (113) to the d’Alembert equation $\Phi_{z_{+},z_{-}}$ for $\Phi=\Phi(z_{+},z_{-})$. Thus the generic form of solution for equation (113) is
$M=\frac{1}{\tilde{r}}(F(\tilde{r}+x_{2})+G(\tilde{r}-x_{2}))$ (116)
where $F(\tilde{r}+x_{2})$ and $G(\tilde{r}-x_{2})$ are arbitrary functions whose arguments are fixed in the brackets. The corresponding potential (25) looks as follows:
$V=\frac{W(z_{+})+K(z_{-})}{F(z_{+})+G(z_{-})}.$ (117)
Solving equations (114) where $M$ is function (116) we obtain the related function $N$ in the form
$N={x_{2}}M+G(z_{-})-F(z_{+}).$ (118)
In accordance with the above the PDM system with the mass and potential given by relations (116) and (118) admits the integral of motion of the following form:
$Q=\{L_{3},P_{1}\}+P_{a}x_{2}P_{a}-((F(z_{+})-G(z_{-}))\cdot H)+W(z_{+})-K(z_{-}).$ (119)
The next version of the arbitrary parameters which we consider is given by relations (108) and (110). The corresponding equation (112) is reduced to the following one:
$(x_{2}^{2}-x_{1}^{2})M_{1,2}+x_{1}x_{2}(M_{1,1}-M_{2,2})+3(x_{2}M_{1}-x_{1}M_{2})=0$ (120)
which is solved by the following function:
$M=F(\tilde{r})-\frac{G(\varphi)}{\tilde{r}^{2}}$ (121)
while the related potential $V$ (25) and function $N$ satisfying equations (98) have the following form:
$V=\frac{\tilde{F}(\tilde{r})-\tilde{G}(\varphi)}{F(\tilde{r})-G(\varphi)},\ N=G(\varphi).$ (122)
The PDM system whose mass and potential are given by relations (121) and (122) admits the following integral of motion:
$Q=L_{3}^{2}+(G(\varphi)\cdot H)-\tilde{G}(\varphi).$ (123)
The next version (109), (110) corresponds to the most complicated determining equations. Up to scaling of variables $x_{a}$ we can set $\sigma=\pm 1$ and reduce (96) to the following form:
$\begin{split}&(x_{2}^{2}-1)M_{1}-x_{1}x_{2}M_{2}+N_{1}=0,\\ &(x_{1}^{2}-1)M_{2}-x_{1}x_{2}M_{1}+N_{2}=0,\ N_{3}=0\end{split}$ (124)
while the corresponding differential consequence (112) looks as follows:
$(x_{2}^{2}-x_{1}^{2}-2\sigma)M_{12}+x_{1}x_{2}(M_{11}-M_{22})+3(x_{2}F_{1}-x_{1}F_{2})=0$ (125)
an is rather cumbersome. However, choosing new dependent and independent variables, namely,
$\hat{M}=\omega M,\ \tilde{x}_{1}=\omega+\tilde{r}^{2},\ \tilde{x}_{2}=\omega-\tilde{r}^{2}$ (126)
where $\omega=\sqrt{\tilde{r}^{4}-4y},y=x_{1}^{2}-x_{2}^{2}-\sigma$ we can reduce (112) to the standard D’Alembert equation for $\hat{M}=\hat{M}(\tilde{x}_{1},\tilde{x}_{2})$ whose solutions are
$\hat{M}=F_{1}(\tilde{x}_{1})+F_{2}(\tilde{x}_{2})$ (127)
where $F_{1}(.)$ and $F_{2}(.)$ are arbitrary functions. Thus the generic solution for the mass function is:
$M=\frac{F_{1}(\tilde{x}_{1})+F_{2}(\tilde{x}_{2})}{\omega}.$ (128)
For any fixed functions $F_{1}(.)$ and $F_{2}(.)$ we can solve equation (124) and find the corresponding function $N$. Unfortunately it is impossible to find $N$ in closed form for $F_{1}(.)$ and $F_{2}(.)$ arbitrary. However, it can be done for the special class of these functions. Namely, considering the polynomial functions
$F_{1}(\tilde{x}_{1})+F_{2}(\tilde{x}_{2})=c_{m}\Omega_{m}$ (129)
where
$\Omega_{m}={(\omega+\tilde{r}^{2})^{m}+(-1)^{m+1}(\omega-\tilde{r}^{2})^{m}}$ (130)
and summation is imposed over the repeating indices over all natural values $m=1,2,..$
Substituting the corresponding expression for $M$, i.e. $M=\frac{c_{m}\Omega_{m}}{\omega}$ into equations (124) we come to the relatively simple solvable system whose solutions are
$N=\frac{-2yc_{m}\Omega_{m-1}}{\omega}.$ (131)
At this point we stop our discussion of the most complicated version (109), (110) Notice that the corresponding integrals of motion (96) is reduced to the following form
$Q=L_{3}^{2}+(P_{1}^{2}-P_{2}^{2})+\left(N(F,G)\cdot H\right)-\tilde{N}(\tilde{F},\tilde{G})$ (132)
were $N(F,G)$ and $\tilde{N}(F,G)$ are functions specified in the above.
The last version we are supposed to consider is presented by relations (111). The corresponding equation (112) is reduced to the standard D’Alembert form $M_{1,2}=0$ and is solved by the following function:
$M=F(x_{1})+G(x_{2}).$ (133)
The related potential $V$ (25) and function $N$ satisfying condition (98) have the following form:
$V=\frac{\tilde{F}(x_{1})+\tilde{G}(x_{2})}{F(x_{1})+G(x_{2})},\ N_{(1)}=G(x_{2})-F(x_{1})$ (134)
and the corresponding integral of motion (96) is reduced to the following form:
$Q=P_{1}^{2}+\left((G(x_{2})-F(x_{1}))\cdot H\right)+\tilde{F}_{1}(x_{1})-\tilde{G}(x_{2}).$ (135)
Our search for shift invariant PDM systems admitting a second order integral of motion in form (97) turned to the end. The classification results obtained in this and previous sections are summarized in Table 2.
4.2 Search for superintegrable systems
In previous subsection we classified integrable system which admit at least one second order integral of motion. The found masses and potentials are defined op to arbitrary functions.
For some special forms of these functions the system can obtain more integrals of motion and be superintegrable or even maximally superintegrable. Just just systems are classified in the present section.
There exist three versions of the considered superintegrable systems. Namely, they can admit two integrals of motion of type (96), two integrals of motion of type (96), or integrals of motion of the both types. All these versions are studied in the following subsections.
4.2.1 Syperintegrable systems admitting integrals of motion of type (96)
The PDM systems admitting one integral of motion of type (96) are classified in section 3.2.1 where four inequivalent systems having this property are presented. The related integrals of motion are given by equations (119), (123), (132) and (135). We are supposed to consider the cases when these integrals of motion are admitted together with one more integral of motion of generic form (96). Fortunately, the cases when such defined pairs of integrals of motion include (123) or (132) can be omitted as being equivalent to the remaining cases up to shift and rotation transformation, and it is sufficient to restrict ourselves to the cases (119) and (135).
Let us start with the system whose mass depends on two arbitrary functions an is presented by equation (133). Such system admits integral of motion (135). Our task is to search the cases when such system admits one more integral of motion of the generic form (96) which can happen for some special forms of the mentioned functions.
Substituting (123) into equation (112) were without loss of generality we can set $\omega=0$ we obtain the following relation:
$((\sigma x_{2}-\nu)x_{1}-\kappa)(F_{1,1}-G_{2,2}+3(\sigma x_{2}-\nu)F_{1}-3\sigma x_{1}G_{2}=0.$ (136)
Let $\sigma\neq 0$ then we can set $\sigma=1$ and nullify $\nu$ making the shift $x_{2}\to x_{2}+{\nu}$. As a result we come to equation (136) which is solved by the following functions:
$F(x_{1})=c_{1}x_{1}^{2}+c_{2},\ G(x_{2})=c_{1}x_{2}^{2}+c_{3}.$
The corresponding function (133) has the additional Lie symmetry with respect to rotations around the third coordinate axis and so can be ignored.
If in addition $\kappa=0$ equation (136) is solved by functions $F(x_{1})=c_{1}x_{1}^{2}+c_{2}x_{1}^{-2},\ G(x_{2})=c_{1}x_{2}^{2}+c_{3}x_{2}^{-2}+c_{4}$. The related functions $M$ (133) and $N$ look as follows:
$M=c_{1}\tilde{r}^{2}+\frac{c_{2}}{x_{1}^{2}}+\frac{c_{3}}{x_{2}^{2}}+c_{4},\ N=-\frac{c_{2}x_{2}^{2}}{x_{1}^{2}}-\frac{c_{3}x_{1}^{2}}{x_{2}^{2}}$ (137)
and integral of motion (96) is reduced to the following form:
$Q=L_{3}^{2}-(N\cdot H)+\frac{c_{6}x_{2}^{2}}{x_{1}^{2}}+\frac{c_{7}x_{1}^{2}}{x_{2}^{2}}.$ (138)
If parameter $\sigma$ in (136) is trivial but $\nu\neq 0$ we set $\nu=1$ and nullify $\kappa$ making the shift $x_{1}\to x_{1}-{\kappa}$. Solutions of the such reduced equation (136) generate the following function $M$ of generic form (133), potential $V$ (25) and function $N$ solving equations (98):
$\begin{array}[]{l}M=\frac{c_{1}}{x_{1}^{2}}+c_{2}(x_{1}^{2}+4x_{2}^{2})+c_{3}x_{2}+c_{4},\\ V=\frac{c_{1}+c_{2}(x_{1}^{2}+4x_{2}^{2})+x_{1}^{2}(c_{3}x_{2}+4_{8})}{c_{5}+c_{6}(x_{1}^{2}+4x_{2}^{2})+x_{1}^{2}(c_{7}x_{2}+c_{8})},\ \ N=\frac{1}{2x_{1}^{2}}(4x_{2}(c_{1}-c_{2}x_{1}^{4})-c_{3}x_{1}^{4}).\end{array}$ (141)
Finally, if both $\sigma$ and $\nu$ are trivial, equations (136) and (98) are easy solvable and generate the following $M,V$ and $N$:
$\begin{array}[]{l}M=c_{1}x_{1}+c_{2}x_{2}+c_{3}\tilde{r}^{2}+c_{4},\\ N=-(2c_{3}x_{1}x_{2}+c_{2}x_{1}+c_{1}x_{2}),\\ V=\frac{c_{5}x_{1}+c_{6}x_{2}+c_{7}\tilde{r}^{2}+c_{8}}{c_{1}x_{1}+c_{2}x_{2}+c_{3}\tilde{r}^{2}+c_{4}}.\end{array}$ (145)
The PDM systems whose mass and potential are fixed by relations (141) and (145) admit the second integrals of motion in form (96) where functions $\tilde{N}$ has the same form as $N$ where, however, parameters $c_{k}$ are changed to $c_{k+4}$, and values of parameters $\sigma,\nu,\kappa$ are specified in the above.
The next step is to consider the system admitting integral of motion (123). The corresponding mass (121) satisfies equation (120) and depends on two arbitrary function. Out task is to find such special form of these functions corresponding to the case when such system admits one more integral of motion of the generic form (96). Substituting It happens in the case when the system of equations (113) and (108) is consistent.
Substituting (121) into (108) we come to the following equation for functions $F(\tilde{r})$ and $G(\varphi)$:
$\begin{array}[]{l}\left(\kappa\cos(2\varphi)-\frac{\omega}{2}\sin(2\varphi)\right)(\tilde{r}^{5}F_{\tilde{r},\tilde{r}}-\tilde{r}^{4}F_{\tilde{r}}+2\tilde{r}G_{\varphi,\varphi}-8G_{\varphi})\\ +\nu\cos(\varphi)(\tilde{r}^{6}F_{\tilde{r},\tilde{r}}+4\tilde{r}^{5}F_{\tilde{r}}+\tilde{r}^{2}G_{\varphi,\varphi}-2G)+3\sin(\varphi)\tilde{r}^{2}G_{\varphi}=0\end{array}$ (148)
where by definition $\nu\neq 0$ since for $\nu$ zero we come to the case rotationally equivalent to (141).
Equation (148) is consistent for arbitrary values of parameters $\nu,\kappa$ and $\omega$. However, if $\kappa$ is not trivial, its solutions are $F(r)=c_{1}+\frac{c_{2}}{\tilde{r}^{2}}$ and $G(\varphi)=c_{2}$ , and the related mass (121) is constant.
Let both $\kappa$ and $\omega$ be trivial then equation (148) is solved by the following function:
$F(\tilde{r})=c_{1}+\frac{c_{2}}{\tilde{r}}+\frac{c_{5}}{\tilde{r}^{2}},\ G(\varphi)=c_{5}-\frac{c_{3}\sin(\varphi)}{\cos(\varphi)^{2}}-\frac{c_{4}}{\cos(\varphi)^{2}}$ (149)
while the corresponding mass (121), potential (25) and function $N$ solving equations (98) are:
$\begin{split}&M=c_{1}+\frac{c_{2}}{\tilde{r}}+\frac{c_{3}}{x_{1}^{2}}+\frac{c_{4}x_{2}}{\tilde{r}x_{1}^{2}},\\ &V=\frac{c_{6}x_{1}^{2}+c_{8}x_{2}+c_{5}\tilde{r}x_{1}^{2}+c_{7}\tilde{r}}{c_{2}x_{1}^{2}+c_{4}x_{2}+c_{1}\tilde{r}x_{1}^{2}+c_{3}\tilde{r}}.\end{split}$ (150)
and
$N=x_{2}\left(\frac{c_{2}}{\tilde{r}}+\frac{c_{3}}{x_{1}^{2}}\right)+\frac{c_{4}(x_{1}^{2}+2x_{2}^{2})}{\tilde{r}x_{1}^{2}}.$ (151)
The related PDM system admits two second order integrals of motion, namely, operator (123) with $G(\varphi)$ given by relation (149) where $c_{5}=0$ and operator given by relation (96) where the only nonzero parameter is $\nu=1$ and $N$ is function (151).
If $\kappa=0$ but both $\nu$ and $\omega$ are nontrivial, equation (148) is solved by functions (149) with $c_{2}=c_{4}=0$. The corresponding PDM system admits two the above mentioned integrals of motion with trivial $c_{2}$ and $c_{3}\neq 0$ and one more integral of motion $Q=P_{1}^{2}-\left(\frac{c_{3}}{x_{1}^{2}}\cdot H\right)+\frac{c_{6}}{x_{1}^{2}}$. The obtained results are presented in Item 3 of Table 4.
4.2.2 Superintegrable systems admitting integrals of motion of type (97)
Consider now superintegrable systems admitting two integrals of motion of generic form (97). The rules of game are analogous to ones used in the previous subsection. Namely, to classify such systems it is necessary to consider step by step all inequivalent versions of mass functions presented in section 4.1 and test their compatibility with generic equation (101).
Let us start with the mass function presented in (105). Substituting it into (101) where $\lambda\to\tilde{\lambda}$ we reduce the latter equation to the following form:
$(\alpha x_{2}+\mu x_{1}+\lambda^{32})F_{2}+2\alpha F-\lambda\mu x_{2}+3\alpha\lambda x_{1}+\tilde{\lambda}=0$ (152)
where $F=F(x_{2})$.
Equation (152) has consistent solutions for $F$ in the unique case, namely, when $\mu=\alpha=0,\lambda^{32}\neq 0$. Choosing such values of the arbitrary coefficients and setting $\lambda^{32}=1$ we come to the following solution:
$F=-\tilde{\lambda}x_{2}+c_{1}$ (153)
which generates the following function (105) and potential (25):
$\begin{split}&M=\lambda x_{1}-\tilde{\lambda}x_{2}+c_{1},\\ &V=\frac{c_{1}x_{1}+c_{3}x_{2}+c_{3}}{\lambda x_{1}-\tilde{\lambda}x_{2}+c_{1}}.\end{split}$ (154)
The PDM system with the mass and potential presented in (154) admits two second order integrals of motion of type (97), namely
$Q_{1}=\{P_{3},P_{1}\}+\lambda x_{3}$ (155)
and
$Q_{2}=\{P_{3},P_{2}\}+\tilde{\lambda}x_{3}.$ (156)
Moreover, it admits additional integrals of motion as it will be shown in the next subsection.
4.2.3 Superintegrable systems admitting integrals of motion of both types
Let arbitrary parameters satisfy conditions (102). The related function solves equation (96) and $M$ is defined up to arbitrary function which depend on a single variable, and our task is to specify the form of this function and admissible values of arbitrary parameters which are compatible with equation (112).
In the case when $M$ has the form presented in (102) equation (112) is reduced to the ordinary second order equation for function $F=F(z)$ where $z=\alpha\ln(\tilde{r})-\mu\varphi$. It has the following form
$\Phi_{(2)}F_{z,z}+\Phi_{(1)}F_{z}+\Phi_{(0)}F+\Phi=0$ (157)
where $F_{z}=\frac{\partial F}{\partial z}$, etc., $\Phi_{(0)}=2\nu x_{1}\tilde{r}^{2}-8(\omega x_{1}x_{2}+\kappa(x_{2}^{2}-x_{1}^{2})),\Phi=0,$ $\Phi_{(2)}$ and $\Phi_{(1)}$ are fourth order polynomials in $x_{1},x_{2}$.
Since $\Phi_{(0)}$ is not a function of $z$ it follows from (157) that either $\Phi_{(0)}$ or $F$ should be trivial. To obtain a nontrivial $F$ we have to nullify $\Phi_{(0)}$ which can be achieved by setting $\omega=\kappa=\nu=0$. As a result equation (112) is reduced to the following form:
$0\displaystyle 2\sigma\mu F_{z,z}=$ 0 . (158)
Since $\mu$ by definition is nontrivial it follows from (158) that there are two possibilities: $\sigma=0$, $F(z)$ is an arbitrary function, and
$F=c_{1}z+c_{2},\sigma\neq 0.$ (159)
In the first case we have generic solution (102) for $M$, and the only integral of motion is given by relation (96) where the only nonzero coefficients are $\mu$ and $\alpha$, i.e.,
$Q^{(1)}=\mu P_{3}L_{3}+\alpha\{P_{3},D\}-\left(2c_{2}\nu x_{3}\cdot H\right)+2c_{5}\nu x_{3}.$ (160)
In the case (159) there are two integrals of motion, namely, (160) and
$Q^{(2)}=L_{3}^{2}+2(c_{1}\varphi\cdot H)-2c_{4}\varphi.$ (161)
The latter integral of motion is the particular case of (98) where the only nonzero coefficient is $\sigma$.
The next version we consider is presented in equation(103). Substituting the function $M$ given there into equation (112) we again come to equation of generic form (157) where $z=\tilde{r}$ and
$\begin{split}&\Phi_{(2)}=-4\mu\tilde{r}^{4}(\nu x_{1}\tilde{r}^{2}-\omega x_{1}x_{2}+\kappa(x_{1}^{2}-x_{2}^{2})),\ \Phi_{(1)}=6\mu\nu x_{1}\tilde{r}^{4},\\ &\Phi_{(0)}=0,\ \Phi=2\lambda\sigma\tilde{r}^{4}-4x_{2}\nu\tilde{r}^{2}+\omega(x_{1}^{2}-x_{2}^{2})+4\kappa x_{1}x_{2}.\end{split}$ (162)
In accordance with (162) equation (157) has nontrivial solutions only in the case when all parameters $\omega,\nu,\kappa$ and $\sigma$ are equal to zero. Thus solution (103) generates the PDM system which admit the only integral of motion in form (96) where the only nonzero coefficient is $\mu$. ,i.e.,
$Q=\{P_{3},L_{3}\}-c_{1}(x_{3}\cdot H)+\lambda x_{3}.$ (163)
Considering version (104) we come to equation (157) where $z=\varphi$ and
$\begin{split}&\Phi_{(2)}=(\nu x_{1}\tilde{r}^{2}+\kappa(x_{1}^{2}-x_{2}^{2})-\omega x_{1}x_{2}),\\ &\Phi_{(1)}=-3\nu x_{2}\tilde{r}^{2}-12\kappa x_{1}x_{2}+3\omega(x_{2}^{2}-x_{1}^{2}),\\ &\Phi=8\omega x_{1}x_{2}-2\nu x_{1}\tilde{r}^{2}+8\kappa(x_{2}^{2}-x_{1}^{2}),\ \Phi_{(0)}=0.\end{split}$ (164)
Functions (164) can be reduced to functions of $\varphi$ in two cases, namely, when coefficients $\omega,\kappa,\sigma$ and $\nu$ satisfy one of the following relations:
$\omega=0,\ \kappa=0,\nu\neq 0$ (165)
$\nu=0.$ (166)
Notice that in the second case we can (and have done it) to nullify also coefficient $\kappa$ applying the rotation transformation.
In case (165) the corresponding equation (157) is solved by the following function:
$F(\varphi)=\frac{c_{1}}{\cos(\varphi)^{2}}+\frac{c_{2}\sin(\varphi)}{\cos(\varphi)^{2}}$ (167)
while for the case (166) we obtain
$F(\varphi)=\frac{c_{1}}{\cos(\varphi)^{2}}+\frac{c_{2}}{\sin(\varphi)^{2}}.$ (168)
The corresponding masses and potentials look as follows:
$\begin{split}&M=\frac{c_{2}}{x_{1}^{2}}+\frac{c_{1}x_{2}}{\tilde{r}x_{1}^{2}}-{\lambda},\\ &V=\frac{c_{3}\tilde{r}+c_{4}x_{2}+c_{5}\tilde{r}x_{1}^{2}}{c_{1}\tilde{r}+c_{2}x_{2}-\lambda\tilde{r}x_{1}^{2}}\end{split}$ (169)
for case (165), and
$\begin{split}&M=\frac{c_{2}}{x_{1}^{2}}+\frac{c_{2}}{x_{1}^{2}}-{\lambda},\\ &V=\frac{c_{3}x_{1}^{2}+c_{4}x_{2}^{2}+c_{5}x_{1}^{2}x_{2}^{2}}{c_{1}x_{1}^{2}+c_{2}3x_{2}^{2}-\lambda x_{1}^{2}x_{2}^{2}}\end{split}$ (170)
@@ 4.2.3 Superintegrable systems admitting integrals of motion of both types
The PDM systems whose masses and potentials are fixed in (169) and (170) admit the following integrals of motion of type (96)
$Q^{(1)}=\{P_{1},L_{3}\}+\left(\frac{2c_{2}\tilde{r}+c_{1}(x_{1}^{2}+2x_{2}^{2})}{x_{1}^{2}\tilde{r}}\cdot H\right)-\frac{2c_{5}\tilde{r}+c_{4}(x_{1}^{2}+2x_{2}^{2})}{x_{1}^{2}\tilde{r}}$ (171)
and
$P_{1}^{2}-P_{2}^{2}+\left(\frac{c_{1}x_{1}^{2}-c_{2}x_{2}^{2}}{x_{1}^{2}x_{2}^{2}}\cdot H\right)-\frac{c_{4}}{x_{2}^{2}}+\frac{c_{5}}{x_{1}^{2}}$ (172)
correspondingly. In addition, both these systems by construction admit integral of motion (107).
4.2.4 Integrals of motion dependent on $x_{3}$